OpenStudy (anonymous):

Which of the following is equal to 1.5 Liters? 150 mL 1500 cm3 15 m3 15000 mm3

OpenStudy (asib1214):

15m^3

OpenStudy (asib1214):

if cm^3 means ml, m means liters

OpenStudy (asib1214):

M^3

OpenStudy (somy):

L/1000= m3 L*1000= mL= cm3 L*1000000= mm3

OpenStudy (somy):

replace all L by 1.5 and do that small calculation hope it helps))